Solve by complementing the square
1/2x^2 + 3x =17

Answer by mananth(16946)

=

(1/2)x^2 + 3x =1
multiply by 2
x^2+6x =2
Third term can be found out by the formula
(1/2* coefficient of middle term )^2
((1/2)*6)^2
=9

x^2 + 6x+9 =2+9
(x+3)^2 = 11
x+3= +/-sqrt(11)
x= -3 +/- sqrt(11)